OpenStudy (anonymous):

Two mechanics worked on a car. The first mechanic worked for 10 hours, and the second mechanic worked for 15 hours. Together they charged a total of 1825 . What was the rate charged per hour by each mechanic if the sum of the two rates was 160 per hour?

OpenStudy (anonymous):

How would you figure this one out ?

OpenStudy (anonymous):

You need a variable for each mechanic.

OpenStudy (anonymous):

So have \(x\) be the first mechanic's rate, and \(y\) be the second mechanics rate.

OpenStudy (anonymous):

`What was the rate charged per hour by each mechanic if the sum of the two rates was 160` This tells us:\[ x+y=160 \]

OpenStudy (anonymous):

`The first mechanic worked for 10 hours, and the second mechanic worked for 15 hours. Together they charged a total of 1825` This tells us \[ 10x+15y = 1825 \]

OpenStudy (anonymous):

Now we have a system of equations: \[ \begin{split} x&+&y&=&160\\ 10x&+&15y&=&1825 \end{split} \]

OpenStudy (anonymous):

There are two ways we can solve this. First is elimination. Notice the coefficients of the \(x\)s. The first equation has \(1x\) and t he second had \(10x\). We multiply both equations by the co efficient of the other. It's similar to getting like terms in a fraction: \[ 10x+10y=1600\\ 10x+15y = 1825 \]Now we subtract the second equation from the first: \[ -5y = -225 \]

OpenStudy (anonymous):

We solve for \(y\) and then plug that number into our old equation in order to solve for \(x\).

OpenStudy (anonymous):

The second way to solve this is substitution. In this case we isolate a variable in one of the equations:\[ x+y=160 \implies x = 160-y \]Then we substitute this into the other equation: \[ 10x+15y=1825 \implies 10(160-y) + 15y = 1825\implies 5y=225 \]

OpenStudy (anonymous):

Then it follows as it would have with elimination.

OpenStudy (anonymous):

okidoke awesome ! is there any formula i could remember for any other problems like this ?

OpenStudy (anonymous):

and since 5y=225 you would then divide getting your "y" right? then you would go through the problem and solve for "x"?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

There isn't really a general formula.

OpenStudy (anonymous):

Yes, you divide by \(5\) to isolate \(y\).